Operations Report – September 7th

MDRS Complete Operations Report 07-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: Refrigerator in intern trailer

Action is taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power
Solar— SOC % Last 24 hours:
Average 80
Minimum 59
Maximum 100
Note on solar: Nothing to report

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%

Ethanol Free Gasoline – 0 gallons (and it looks like we are missing a couple of gas cans)

Water (loft tank): 50 gallons
Water Meter: n/a units
Water (static tank) – 350 gallons
Static to Loft Pump used – yes
Water in GreenHab – 160 gallons
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 75 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes
Beginning charge: 100
Ending charge:
Currently charging: yes

Notes on rovers: The remaining four rovers, with new batteries, will be delivered to the entrance to Cow Dung Road, at 11 am on Wednesday.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): none
Reason for use: n/a
Oil Added? No
ATV Fuel Used: Not checked Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, hauling water

CrewCar used and why, where? Yes, to Loa for food and supplies

Luna used and why, where? Yes, emptied the remaining materials and put the seat back in

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Radios are charged and are being tested.

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: The tunnels have been completely covered and this year they look really well made as this team really worked on making them secure and looking good.

We completed a bunch of small projects today so we could focus on painting for the remainder of the work party.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: A wooden wall was constructed between the bathroom and bedroom in the intern trailer (I am very jealous as it looks amazing). The bedroom will be reserved for the Assistant Director. The water system was checked and is operational. The septic system was looked at and problems were found that will need to be corrected. The drainage pipes, ditch, and tank were all fixed up and look great.
Summary of GreenHab operations: Water prep still underway

Summary of ScienceDome operations: Tomorrow the lift boom (or whatever it is called) and the power water rentals arrive. We are planning to begin washing the ScienceDome as our main project. Because of the red bentonite clay having an electrical charge, you can’t just wash it off with water. You need to actually physically wash it off with soapy water and then rinse. So basically today we will be washing a building like it was a big baby!

Summary of RAM operations: This building has a lot of peeling paint. We are not going to work on it while the work party is happening, but it will need some attention soon.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Still miserably hot.

Questions, concerns, supplies needed and requests: Nothing at this time.

Copyright © The Mars Society. All rights reserved. | Main Site